OpenStudy (anonymous):

Derivatives. Could someone just tell me what the derivative of 100/x^2 would be?

OpenStudy (jdoe0001):

\(\bf \cfrac{100}{x^2}\implies \cfrac{100}{1}\cdot \cfrac{1}{x^2}\implies 100\cdot x^{-2}\) use the power rule

OpenStudy (anonymous):

okay cool, and then from there does -200/x^3 = the derivative? not sure if i'm following rules there, I've been out of math for a little while